A RECTANGLE IS 16M BY9M.FIND A SIDE OF THE SQUARE WHOSE AREA EQUALS THE AREA OF THE RECTANGLE. BY HOW MUCH DOES THE PERIMETER OF THE RECTANGLE EXCEED THE PERIMETER OF THE SQUARE.
Answer by mananth(16949)

A RECTANGLE IS 16M BY9M.FIND A SIDE OF THE SQUARE WHOSE AREA EQUALS THE AREA OF THE RECTANGLE. BY HOW MUCH DOES THE PERIMETER OF THE RECTANGLE EXCEED THE PERIMETER OF THE SQUARE.
..
Area of rectangle = 16*9= 144 m^2
..
Area of square = L^2 = 144
L= 12 m
..
Perimeter of rectangle = 2*(L+W) ===>2*(16+9)
Perimeter = 50 m
...
Perimeter of square = 4*L = 48 m
..
Difference = 2 m
...
m.ananth@hotmail.ca